Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 20/6/2017

Location : CR21 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • Actions.
    • Hall Manager - water system change-over procedure (stands: the outcome has to be a laminated sheet of instructions in, say, the MLCR for when things go wrong, else the exercise is pointless) - have prescription and draft document (Colin/Josef)
    • Progress on Hall list attachment:MICE-Work-List-May-2017.pdf (Colin) Last remaining item is believed to be verifying the interlocks; hope to get Trevor Hartnett down in next few weeks?
    • CWhyte to request survey of CC Magnets after cycle: need survey of SS's (Tracker positions during data taking) and FC (Absorber location) before pulling the FC out - stands
    • Tracker group: decide what should be done to prevent repeat contamination of Tracker Cryostats
    • Status of miceraid5 disk cleanup (stands) EOverton now has a script but needs to run it: needs to hand over the procedure and script to someone, i.e. Durga stands
    • Investigation of the widespread Watchdog Timer failure
    • Swap of IP numbers of A/C units 2 & 3 - Paolo
  • DC update
  • Hydrogen system update and plans
  • Issues to fix before September : Cycle 2017 Issues : This will be discussed after a prioritisation exercise is carried out.
  • Status of Support for Technical Endeavours And Knowledge (CWhyte)
  • AOB

Minutes

  • Actions
    • Hall Manager - water system change-over procedure (stands) - have prescription and draft document (Colin/Josef), stands
    • Progress on Hall list attachment:MICE-Work-List-May-2017.pdf (Colin) Last remaining item is believed to be verifying the interlocks; hope to get Trevor Hartnett down in next few weeks? No update. Assume Trevor would want to do this an the repaired FC power supply in the same visit?
    • CWhyte to request survey of CC Magnets after cycle: need survey of SS's (Tracker positions during data taking) and FC (Absorber location) before pulling the FC out - FC and SS's surveyed last Thursday - done.
    • Tracker group: decide what should be done to prevent repeat contamination of Tracker Cryostats: Using BIP grade Helium which is unlikley to be contaminated; must ensure we don't accidentally use a lower grade; should test a sample - done
      Still need to decide regarding repeating the warm-up/pump/purge over the summer; have effectively run out of time to do this with Ed in attendance.
    • Status of miceraid5 disk cleanup (stands) EOverton now has a script but needs to run it: needs to hand over the procedure and script to someone, i.e. Durga Is Durga aware of this item?
    • Investigation of the widespread Watchdog Timer failure
    • Swap of IP numbers of A/C units 2 & 3 - Paolo - DNS entries have been swapped; IP numbers swapped in code; AKurup needs to restart the IOC. Stands
      Note that the archived values will continue to be swapped. Note that there should be an obvious discontinuity when viewing the history. Make an eLog entry!
  • DC update
    Mark has measured more piping. Will remove FC turret after inspection of the new one.
    The FC has been surveyed and pulled offline.
  • Hydrogen system update and plans
    R9: The pipework on the R9 turret is almost done. The heat sinking is done. The heat shield will be done Thu-Sat. - can then swap turrets. Need to re-test connections.
    Must measure existing heater element properties and cross-check with the old/Hall turret.
  • Issues to fix before September : Cycle 2017 Issues : This will be discussed after a prioritisation exercise is carried out.
    AKurup and HNebrensky will work through the list in the next couple of days and identify C&M actions. Remaining list will be need to be actioned somehow.
  • Status of Support for Technical Endeavours And Knowledge (CWhyte)
    Yes!
  • AOB
  • Repaired FC PSU: THartnett has been chasing Thurlby-Thandar, no useful outcome.

Updated by Nebrensky, Henry about 5 years ago · 13 revisions